Circular No. 7008

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1550–564 D. A. Smith and R. Remillard, Massachusetts Institute of Technology

(MIT), on behalf of the RXTE/ASM teams at MIT and Goddard Space Flight Center (GSFC), reports: “A transient x-ray source has been detected with the RXTE/ASM near α = 15^{h}50^{m}41^{s}, δ = −56^{o}27′.6 (equinox 2000.0). The source was first detected at an intensity of ∼ 70 mCrab (2–12 keV; 5σ significance) in an observation at Sept. 7.0927 UT. At Sept. 7.4282, the source was observed to have a strength of 170 ± 20 mCrab. Crossed lines of position were obtained; from these, the source is estimated to be located within a region (90-percent confidence) defined by a box of width 4′.3 and length 11′.0 with corners at α = 15^{h}51^{m}31^{s}, δ = −56^{o}33′.4; 15^{h}50^{m}19^{s}, −56^{o}26′.1; 15^{h}49^{m}52^{s}, −56^{o}19′.3; 15^{h}50^{m}57^{s}, −56^{o}25′.1.”

F. E. Marshall, GSFC; and E. A. Smith, Hughes STX and GSFC, write: “An RXTE PCA observation on Sept. 7.72 UT of the transient XTE J1550–564 shows a highly-variable source with an average 2–10-keV flux of 6 × 10^{−8} erg s^{−1} cm^{−2}. Aperiodic variability of about a factor of two was seen on timescales of less than a second. The source is likely to be a black-hole transient.”

XTE J1906+09 T. Takeshima and R. H. D. Corbet, Goddard Space Flight Center

(GSFC) and Universities Space Research Association; and J. H. Swank, GSFC, on behalf of the RXTE Science Operations Center, report: “During a series of pointed observations of SGR 1900+14 with the Proportional Counter Array on RXTE, a coherent pulsation at period of 89.178 ± 0.003 s (not barycenter-corrected) was detected, in addition to 5.16-s pulsations (IAUC 7001, 7002, 7003, 7004, 7005), in data obtained on Aug. 31.34– 31.52, Sept. 1.24–1.52, and 2.41–2.52 UT. However, the 89-s pulsations were not seen in observations on Aug. 29.41–29.59 and 30.28–30.46. This period and the 2–10-keV pulse profile are consistent with those reported by Marsden et al. (1998, Ap.J. 502, L129) for XTE J1906+09. The average pulsed PCA count rates were 1.6, 2.5, and 4.3 counts/s (2–10 keV) in the three observations, respectively. As SGR 1900+14 pulses at a 5.16-s period, XTE J1906+09 with its 89-s period is probably a transient pulsar and not associated with the soft γ-ray repeater itself. Offset pointed observations with the PCA are scheduled to attempt to determine the source position.”

1998 September 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green